Status Report

SOHO Daily Meeting Minutes for Monday, 17 June 2002

By SpaceRef Editor
June 17, 2002
Filed under , ,

* RHESSI target for today is AR 0001; target for tomorrow
is TBD.

* Any team that has constraints for the August 20 offpoint,
please let the SOC know by the end of June.

* Security scans will probably be postponed from Wednesday to
a later date.

* PDU Maintenance on June 20 at 12:40pm to 3:20pm local time.

Spacecraft Status: Nominal

Spacecraft Anomalies:
167/1743 Single DMA Failure Over One Format.

Accomplished Activities:
DOY 165: VIRGO, SWAN, RSL, New control star, mag = 4.65

Planned Activities: VIRGO, SWAN

Upcoming Operations: None

Ground Anomalies:
164/1830 SH5WS1 rebooted itself, cause unknown. String 5
was functioning as an offline string at the time.
165/1516 Wayward command delayed NRT activities 2.5 minutes.
165/2025 Wayward command delayed NRT activities 9 minutes.
165/2036 Wayward command delayed NRT activities 12 minutes.
165/2103 Wayward command delayed NRT activities 11 minutes.
166/1648 Wayward command delayed SVM commanding 3 minutes.
166/0121 D46 unable to deliver LMC data for 39 minutes,
cause unknown.
167/0800 Invalid clock correlation data due to a TGC failure
at D66.
167/1849 Wayward command delayed NRT activities 19 minutes.
167/1918 Wayward command delayed NRT activities 4 minutes.
168/0011 Lost command ability for 4 minutes 47 seconds due
to a CPA failure at D27. Delayed end of pass
168/0430 Received degraded TLM from 0435 until 0508 (33
minutes), cause unknown. Lost data should be
recoverable from the CDR.

06/14 DOY 165 – 06/17 DOY 168

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. JOP146 and polarity reversal study. Tomorrow: full
sun observations and observing AR0001.

UVCS: Nominal. Coronal hole observations and mini-synoptic.
Tomorrow: Cassini conjunction observations and synoptic.

LASCO: Nominal. C2/C3 synoptics, with higher cadence during EIT

EIT: Nominal. Bakeout until Tuesday morning, resuming 195 CME
watch early Wednesday.

MDI: Nominal. Full disk magnetograms and dopplergrams until
this afternoon then high resolution observations.

TRACE: Nominal. Full disk mosaic in 171 and observing AR 9998 and AR 9997.

SpaceRef staff editor.